OpenStudy (anonymous):

If f(X)= 5-2x, find f^(-1)(3). The book answer is 1, how did they get that?

OpenStudy (anonymous):

I'll have a go: let x= 5-2y 2y = 5-x y = (5-x)/2 So f(x)^-1 = (5-x)/2 f(3)^-1 = (5-3)/2 =2/2 = 1 EZ Before you go can you tell me what the recommended notation is for inverting functions in your text? I always use y=x but you may have something different.

OpenStudy (anonymous):

I use y=x as well. I hope that answers your question. BTW thanks so much for helping me!